As a person moves about in a dry environment, electric charge accumulates on the person's body. Once it is at high voltage, either positive or negative, the body can discharge via sparks and shocks. Consider a human body isolated from ground, with the typical capacitance 130 pF. (a) What charge on the body will produce a potential of 15.0 kV? (b) Sensitive electronic devices can be destroyed by electrostatic discharge from a person. A particular device can be destroyed by a discharge releasing an energy of 240 µJ. To what voltage on the body does this situation correspond?
Answer #1

Using the formula Q = CV, the charge is...

Q = (130 x 10^-12 F) x (15.0 x 10^3 V) = 1.95 µC

(b) In this situation, we will use the formula U = .5CV^2, and rearrange that into V =sqrt 2U / C to obtain the voltage.

V = sqrt(2 x (240 x 10^-6 J)) / (130 x 10^-12 C) = 1923.5 V